OpenStudy (anonymous):

2. Determine if {-2/3 , 5} is a solution to -3x + 2y = -4. Please show all of your work

OpenStudy (anonymous):

this is means that you need to substitute x=-2/3 and y=5 in your equation & verify if the stated is true. try it...

OpenStudy (anonymous):

-3 * (-2/3) +2*5=-4 ...?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

did you get the answer? what is it?

OpenStudy (anonymous):

2+10=-4 which is not true. so, {-2/3 , 5} is NOT a solution, right?

OpenStudy (anonymous):

yes that's what i got your right